OpenStudy (brydenwright):

Find all solutions to the equation. cos x = - 1/2

OpenStudy (kg1975):

x1 = pi/3 + k.2 pi x2 =-pi/3 + k.2 p

OpenStudy (kg1975):

If cos x = 1/2 then x = cos^(-1) (1/2) therefore, x = 60 degrees , 300 degrees {where x is greater than or equal to 0 and lesser than or equal to 360} or pi/3 radians, 5pi/3 radians {where x is greater than or equal to 0 and lesser than or equal to 2pi}

OpenStudy (kg1975):

Does that make sense? To You?

OpenStudy (brydenwright):

These are my answer choices

jhonyy9 (jhonyy9):

hope you know that cos 60° = 1/2 so now just you need to know where cosine has negative values ,in what quadrant hope helped

OpenStudy (kg1975):

With My Explanation before the answer is C sometimes ppl dont understand so i try my best to explain it but if they dont get it then i give the answer